# STANCE-$\text{C}^{3}$: Domain-adaptive Cross-target Stance Detection via Contrastive Learning and Counterfactual Generation

### v1

@@ VI Conclusion
The robustness of NLP models across domains and multiple targets of interest remains an open area of work in stance detection tasks. In this paper, we proposed a stance detection model STANCE-$\text{C}^{3}$ that exhibited better performance over existing SOTA for cross-domain, multi-target scenarios. By conducting experiments using datasets from the COVID pandemic, we showed that cross-domain and cross-target performance generally degrades for all models. However, in comparison to models such as MoLE and MTL, the cross-domain degradation performance of STANCE-$\text{C}^{3}$ is significantly lower, providing some evidence of its robustness across domains. Additionally, STANCE-$\text{C}^{3}$ generally outperforms the SOTA for multiple, cross-domain targets. Ablation studies on STANCE-$\text{C}^{3}$’s novel components namely, the contrastive loss function and the use of domain counterfactuals for data augmentation, illustrate the marginal impacts of these components on performance and explain why STANCE-$\text{C}^{3}$performed better than its SOTA counterparts. Future work includes extending the proposed training strategy to parameter-efficient models on large language models, along with application to the zero-shot stance detection tasks and other benchmark datasets.
Acknowledgments
This material is based on work supported by the U. S. Department of Homeland Security under Grant Award Number 17STQAC00001-05-00222Disclaimer: ”The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of Homeland Security.”.

### v2

@@ VI Conclusion
The effective use of social media data in facilitating effective public discourse and policy interventions requires agile and adaptable NLP models. Stance detection, as an NLP task, could be a potent tool in accurately gauging public and group-level support or oppositions for critical societal topics. The primary challenge addressed in this work is the difficulty in adapting a pre-trained model to different domains and targets of interest, which limits its utility in time-sensitive applications, such as using social media data to determine public opinions on non-pharmaceutical interventions during a pandemic. Our proposed approach uses counterfactual data generation for extended contexts in which data could be sparse. We showed through ablation studies that this approach improves stance detection across different domains and new targets. Further, we showed that using a modified contrastive loss function complements the data augmentation strategy with much less training data required for the target domain. The proposed architecture in this paper, which combines the two strategies, has shown evidence of outperforming existing SOTA methods in COVID datasets. For applications wherein specialized NLP models present distinct advantages over general purpose models, we showed that aggregating marginal gains by combining strategies is a step in the right direction for robust and adaptable stance detection.